Circular No. 8880

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 T4 (GIBBS) A. R. Gibbs reports his discovery of a comet on unfiltered CCD im-

ages obtained with the 0.68-m Schmidt telescope at Catalina (discovery observation tabulated below), with four co-added 30-s exposures showing a compact coma of size 8′′.5 × 6′′ with a tapered 25′′ tail in p.a. 305^{o}. Four 30-s images taken by R. A. Kowalski with the Mt. Lemmon 1.5-m reflector on Oct. 14.5 UT show a coma of diameter ≈ 10′′ with a straight tail ≈ 25′′ long in p.a. 300^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, J. W. Young writes that his CCD images taken with the Table Mountain 0.61-m telescope on Oct. 15.5 show a round, 6′′-diameter, diffuse coma with a small bright central core and a 30′′ tail in p.a. 290^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Oct. 12.46227  9^{h}57^{m}24.^{s}47  +34^{o}12′33′′.6  18.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-T113.

T = 2007 Aug. 11.396 TT  ω = 49.^{o}967
Ω = 41.159 }2000.0
q = 2.19536 AU  i = 24.007

COMET C/2007 T5 (GIBBS) Gibbs also reports his discovery of another comet with a diffuse 15′′

coma and a brighter 6′′-diameter center (no visible tail) on four co-added 30s unfiltered Catalina Sky Survey exposures (discovery observation tabulated below). Kowalski adds that four co-added 60-s exposures at Mt. Lemmon on Oct. 14.5 UT show a strongly condensed coma of diameter 30′′; his four co-added 75-s exposures on Oct. 15.5 show a condensed coma ≈ 12′′ across and an extended coma ≈ 25′′ across with no visible tail in poor seeing. Young found a very faint diffuse round coma of diameter 26′′ with no tail and very little central condensation on his images from Oct. 15.5.

2007 UT  α_{2000}  δ_{2000}  Mag.
Oct. 13.45523  8^{h}16^{m}28.^{s}09  +19^{o}11′10′′.5  18.2

The available astrometry, preliminary parabolic orbital elements (T = 2008 Feb. 13.463 TT, q = 4.40844 AU, ω = 17^{o}.415, Ω = 110^{o}.115, i = 44^{o}.126, equinox 2000.0), and an ephemeris appear on MPEC 2007-T114.

2007 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green